OpenStudy (anonymous):

Choose the correct simplification of the expression 3b/a^-2 3a^2b a^2/3b 3a^2/b already simplified

OpenStudy (anonymous):

Use this: \[\large \frac{1}{a^{-2}} = a^2\]

OpenStudy (anonymous):

see 1/a^-x=a^x

OpenStudy (anonymous):

so im confused

OpenStudy (anonymous):

Where??

OpenStudy (anonymous):

i need help on this question

OpenStudy (anonymous):

never mind. 3b/a^-2=3b*(1/a^-2)

OpenStudy (anonymous):

We are giving that.. Use the formula that I have written above..

OpenStudy (anonymous):

now use the formula

OpenStudy (anonymous):